OpenStudy (salazarblack):

can someone help??? How many integer solutions has the inequality:x/x+4<=1/x+1

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{x}{x+4} ≤ \frac{1}{x+1}}\)

OpenStudy (solomonzelman):

cross multiply ••••• \(\Large\color{blue}{ x(x+1) ≤ 1(x+4)}\) \(\Large\color{blue}{ x^2+x ≤ x+4}\) \(\Large\color{blue}{ x^2 ~≤ ~ 4}\) \(\Large\color{blue}{ x^2-4 ~~≤ 0}\) \(\Large\color{blue}{ (x-2)(x+2)≤ 0}\)

OpenStudy (solomonzelman):

can you finish the problem off?

OpenStudy (salazarblack):

i found 2 solutions -2 and 2 ..but i don't know how to find the other 3

OpenStudy (salazarblack):

i got till that point as well..

OpenStudy (solomonzelman):

yes, so now we are getting x ≤ 2 and x≤ -2

OpenStudy (solomonzelman):

how many integers are there that are less than -2 ?

OpenStudy (salazarblack):

-3,-4,-5 and so on XD

OpenStudy (anonymous):

You should have \(-2\le x\le2\) as a solution set, not \(x\le-2\). The zero product property doesn't work for inequalities.

OpenStudy (solomonzelman):

didn't know that zero product prop. doesn't work for.....

OpenStudy (anonymous):

Also, note that \(x=-1\) and \(x=-4\) give undefined expressions, so know to avoid those.

OpenStudy (anonymous):

@SolomonZelman, polynomial inequalities can be tricky. What I do is try to picture it (in this case, a parabola shifted down by 4), and see which interval(s) satisfy the inequality. For \(x^2-4\), the parabola dips below the x-axis between -2 and 2.

OpenStudy (salazarblack):

so then the answers would be -2,-1,0,1 and 2

OpenStudy (salazarblack):

but how do i know if x is<= to some value or x=>.. ?